A. The final grades of students in six subjects where he was enrolled were taken and are shown below.


Subject No of units Final grade

Math 100 3 3.0

Lab science 5 1.5

Soc sci 101 3 2.25

Fil 100 2 2.5

P. E 1 3 2.75

NSTP 1 6 1.75


1. Determine his weighted mean grade.

2. If the subjects were equal to the number of units, what would college was average?




1
Expert's answer
2021-12-20T16:27:43-0500

1.

mean = "\sum x_if_i/n=\frac{3\cdot3+5\cdot1.5+3\cdot2.25+2\cdot2.5+3\cdot2.75+6\cdot1.75}{22}=2.13"


2.

mean = "\frac{3+1.5+2.25+2.5+2.75+1.75}{6}=2.29"